Assume that a sample is used to estimate a population mean μ. Find the margin of error M.E. that corresponds to a sample of size 24 with a mean of 51.6 and a standard deviation of 10.5 at a confidence level of 99%. Report M.E. accurate to one decimal place because the sample statistics are presented with this accuracy. Answer should be obtained without any preliminary rounding. However, the critical value may be rounded to 3 decimal places.
Answer #1

Solution:

margin of error

=E=t crit*s/sqrt(n)

alpha=0.01

alpha/2=0.01/2=0.005

degrees of freedom=n-1=24-1=23

t crit in excel==T.INV(0.005;23)=2.80734

so E=2.80734*10.5/sqrt(24)

=6.01698

=6.0

Alternatively with excl formula, margin of error=E

=CONFIDENCE.T(0.01;10.5;24)

=6.016972467

ANSWER:

E=6.0